OpenStudy (anonymous):

Subtract and write the answer in simplest form: 9x over quantity 3x minus 4 minus 12 over quantity 3x minus 4

OpenStudy (anonymous):

\[\frac{9x}{3x-4}-\frac{12}{2x-4}=\frac{9x-12}{3x-4}\]